HF PROMETI Homework problem & discussion questions 1 and 3. Also do problem number 1 Two projects are being considered. The first is estimated to have a $100,000 return each year in todays money and the second is estimated to have a $75,000 return. However, the first is four year while the second is a 6 year. Discount rate is 20% for both. Only one can move forward. Make a decision.
Answer #1

Present value of the first project = 100000*(1-1/(1+R)^n)/R

Present value of the first project = 100000*(1-1/1.20^4)/.2

Present value of the first project = $258873.5

Present value of the Second project = 100000*(1-1/(1+R)^n)/R

Present value of the second project = 75000*(1-1/1.20^6)/.2

Present value of the second project = $249413.3

Here, the decision should be to select the first project, because it has relatively higher present value in today's dollars.